# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ Abstract.
We prove that every irreducible component of the Kollár-Shepherd-Barron and Alexeev (KSBA) moduli space of stable log Calabi–Yau surfaces is a finite quotient of a projective toric variety. This verifies a conjecture of Hacking–Keel–Yu. The proof combines tools from log smooth deformation theory, the minimal model program, punctured log Gromov–Witten theory and mirror symmetry.
@@ 1.1. Context and overview
The higher dimensional analogue of a stable curve, introduced by Kollár–Shepherd-Barron [65] and Alexeev [5], is a stable pair $(Y,B)$, consisting of a projective variety $Y$ and a $\mathbb{Q}$-divisor $B$ such that $(Y,B)$ has semi-log-canonical (slc) singularities, and satisfying the stability condition requiring $K_{Y}+B$ to be $\mathbb{Q}$-Cartier and ample – see [66] for a comprehensive treatment of the KSBA theory. The geometry of the KSBA moduli space of stable pairs has been investigated so far in a handful of cases, including toric and abelian varieties [6], K3 surfaces [7, 8, 10, 71], elliptic surfaces [20, 21], and for some surfaces of general type [9]. Most relevant to the present paper, Alexeev showed in [6] that for toric varieties the normalization of the KSBA moduli space is toric, with associated fan the secondary fan defined as in [37].
More generally, for maximal log Calabi–Yau varieties, it has been conjectured by Keel [62] around 2012 that the KSBA moduli space shall still admit a finite cover by a toric variety. Furthermore, Hacking–Keel–Yu conjectured that such a toric variety should be defined by a generalized version of the “secondary fan” coming from mirror symmetry [49]. We prove this conjecture in dimension two, that is, for all polarized maximal log Calabi–Yau surfaces $(Y,D,L)$ consisting of a projective surface $Y$, with a reduced, anticanonical divisor $D\subset Y$, such that $(Y,D)$ has at worst log canonical singularities, and an ample line bundle $L$ on $Y$. Furthermore, $(Y,D)$ is required to be maximal, in the sense that $D$ is non-empty and singular.
The KSBA moduli space for $(Y,D,L)$, denoted by $\mathcal{M}_{(Y,D,L)}$, is the closure in the moduli space of stable pairs of the stable pairs deformation equivalent to pairs of the form $(Y,D+\epsilon\,C)$, where $C\in|L|$ and $10<\epsilon<\!\!<$ 1. To investigate the geometry of $\mathcal{M}_{(Y,D,L)}$ via mirror symmetry, we first construct a “semistable mirror” to $(Y,D,L)$, given by an open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\to\Delta=\mathrm{Spf}\,\mathbb{C}\llbracket t\rrbracket$, obtained as a projective crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}_{\mathscr{P}}$ of a formal affine canonical 3-fold singularity $\overline{\mathcal{X}}_{\mathscr{P}}$. We then define the “secondary fan” $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, as a coarsening of the Mori fan for the semistable mirror to $(Y,D,L)$. Our main result, Theorem 8.11, states:
@@
The KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface $(Y,D,L)$ admits a finite cover
$S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\longrightarrow\mathcal{M}_{(Y,D,L)}\,,$
@@ 1.2. Motivation from mirror symmetry
Given a stable log Calabi–Yau surface $(Y,D+\epsilon\,C)$ with an ample line bundle $L$, let $U:=Y\setminus D$, and $s$ be the section of $L|_{U}$ defining $C|_{U}$. One can naturally associate to $(Y,D+\epsilon\,C)$ the non-compact Calabi–Yau 3-fold $Z$ given by the hypersurface with equation $xy=s$ in the total space of the rank two vector bundle $L|_{U}\oplus\mathcal{O}_{U}$ over $U$, where $x$ and $y$ are linear coordinates along the fibers. Said differently, $Z$ is the affine conic bundle over the non-compact Calabi–Yau surface $U$ branched over the curve $C|_{U}$. Heuristically, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ can be viewed as a compactification of a “moduli space of complex structures” on $Z$. Mirror symmetry predicts an identification of $\mathcal{M}_{(Y,D,L)}$ with the “stringy extended Kähler moduli space” of a mirror Calabi–Yau 3-fold $\mathcal{X}$ to $Z$ [76]. More precisely, under mirror symmetry different maximal degenerations of $Z$ should correspond to different mirrors related by birational transformations. In this paper, working directly with $(Y,D,L)$ instead of $Z$, we provide a precise realization of this heuristic picture.
Given a choice of maximal degeneration of $(Y,D,L)$ indexed by a choice of polyhedral decomposition $\mathscr{P}$, we give a construction of a “semistable mirror” $\mathcal{X}_{\mathscr{P}}$ as a (formal) non-compact Calabi–Yau 3-fold obtained as a projective crepant resolution of a (formal) canonical 3-fold singularity $\overline{\mathcal{X}}_{\mathscr{P}}$. A mathematically rigorous definition of the “stringy extended Kähler moduli space of $\mathcal{X}_{\mathscr{P}}$” is then provided by the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ with fan given by an explicit coarsening of the Mori fan of $\mathcal{X}_{\mathscr{P}}\rightarrow\overline{\mathcal{X}}_{\mathscr{P}}$.
@@
It is worth noting that in the situation when $Y$ is a toric variety, and $D$ is the toric boundary divisor, mirror symmetry is entirely combinatorial and it essentially reduces to a duality between fans and momentum polytopes: $\overline{\mathcal{X}}$ corresponds to the singular toric variety with associated fan $C(P)$, the cone over the momentum polytope $P$ of $(Y,D,L)$. Choosing a maximal regular subdivision $\mathscr{P}$ and refining $C(P)$ by adding cones over $\mathscr{P}$, one obtains a toric crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}$, the semistable mirror family to $(Y,D,L)$. By standard toric geometry, the Mori fan of this mirror family to $(Y,D,L)$ coincides with the secondary fan of $(Y,D,L)$ [25]. So, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface can be understood combinatorially, either in terms of the secondary fan, or of the Mori fan of the mirror [6, 96]. The existence of a dense torus in $\mathcal{M}_{(Y,D,L)}$ follows in this case from the existence of the basis of $H^{0}(Y,L)$ given by the monomials $z^{m}$ on $(\mathbb{C}^{\star})^{2}=Y\setminus D$ indexed by the integral points $m\in P_{\mathbb{Z}}:=P\cap\mathbb{Z}^{2}$. The key feature of this basis is that, for a curve $C$ defined by an equation $\sum_{m\in P_{\mathbb{Z}}}a_{m}z^{m}=0$ with $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$, the pair $(Y,D+\epsilon\,C)$ is KSBA-stable. Indeed, since for any vertex $m^{\prime}$ of $P$, having $a_{m^{\prime}}\in\mathbb{C}^{\star}$ ensures that $C$ does not intersect the singular point of $D$ corresponding to $m^{\prime}$, it follows that $(Y,D+\epsilon\,C)$ has only slc singularities if $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$ [6].
@@ 1.3. Outline of the paper
In §2, we investigate maximal degenerations of polarized log Calabi–Yau surfaces combinatorially, using real affine geometry. To do so, we consider a “nice” toric model of a polarized log Calabi–Yau $(Y,D,L)$, as in previous work of Engel–Friedman [31], which allows us to systematically describe an associated “Symington polytope” $P$ to $(Y,D,L)$ given by an integral affine manifold with singularities at finitely points [90]. We define the notion of a good polyhedral decomposition on $P$, and we prove the existence of such decompositions in Theorem 2.14. A good polyhedral decomposition defines in particular a maximal degeneration of $(Y,D,L)$, whose central fiber has “intersection complex” $(P,\mathscr{P})$, as explained in §2.4.
To construct a “mirror” to such a maximal degeneration of $(Y,D,L)$, we first construct a normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ with “dual intersection complex” $(P,\mathscr{P})$ in §4.2. The irreducible components of such a normal crossing surface are obtained by $\mathbb{Q}$-Gorenstein smoothings of toric surfaces with quotient Wahl singularities, using Lemmas 4.1 and 4.2, where we show that there are no local to global obstructions for such smoothings. Finally, we appropriately glue these irreducible components to obtain the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$, which has features similar to the central fiber of a Type III Kulikov degeneration in the context of K3 surfaces [35]. A key difference in our context is that the dual intersection complex of $\mathcal{X}_{\mathscr{P},0}$ is a disk, rather than a sphere. So, in §3 we introduce the general notion of an open Kulikov surface, as a normal crossing surface $\mathcal{X}_{0}$ with dual intersection complex a disk, and satisfying some additional properties – see Definition 3.7. After investigating properties of such surfaces, as well as of their affinizations, we show in §4 that for an appropriate gluing the surface $\mathcal{X}_{\mathscr{P},0}$ associated is a $d$-semistable open Kulikov surface. Our key results in Theorems 3.26 and 3.29 prove the following result for a general open Kulikov surface $\mathcal{X}_{0}$:
@@
We refer to such a smoothing $\mathcal{X}\to\Delta$ as an open Kulikov degeneration. By the Minimal Model Program, adapted to the formal set up [74], we conclude in §3.6 that such a Kulikov degeneration is a relative Mori dream space. Hence, we obtain a complete toric fan, the “MoriFan”, in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})\otimes\mathbb{R}$. Then, following [49], we define the “secondary fan” as a “coarsening” of this Mori fan, by identifying neighboring cones connected by “M1 flops” – for details see §6.4, and denote by $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ the corresponding toric variety.
For a polarized log Calabi–Yau surface $(Y,D,L)$ with a choice of $(P,\mathscr{P})$, we refer to the open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ obtained by smoothing $\mathcal{X}_{\mathscr{P},0}$ as the “semistable mirror” of $(Y,D,L)$. We then investigate the “double mirror”, obtained by applying a version of the intrinsic mirror construction of Gross–Siebert [47] to the semistable mirror $\mathcal{X}_{\mathscr{P}}\to\Delta$. One subtlety in our setup is that the semistable mirror is only quasi-projective, while the intrinsic mirror construction of [47] applies to projective degenerations. Therefore, we first prove in Theorem 5.2 that the moduli spaces of punctured log Gromov–Witten invariants of an open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$, defining the structure constants in the intrinsic mirror construction, are proper. The proof crucially uses the fact that we have a projective crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}_{\mathscr{P}}$, where $\overline{\mathcal{X}}_{\mathscr{P}}$ is affine. Hence, analogously as in [47], we define a commutative and associative graded mirror algebra $\mathcal{R}_{\mathcal{X}}$ – see Theorem 5.5. Thus, taking the Proj of this algebra, we obtain an intrinsic mirror family to $\mathcal{X}\to\Delta$ a priori defined over the formal scheme $\mathrm{Spf}\,\mathbb{C}\llbracket NE(\mathcal{X}/\overline{\mathcal{X}})\rrbracket$, as in [47], where $NE(\mathcal{X}/\overline{\mathcal{X}})$ stands for the monoid of integral points in the cone of relative effective curve classes of $\mathcal{X}\rightarrow\overline{\mathcal{X}}$.
@@
By construction [47], the intrinsic mirror algebra $\mathcal{R}_{\mathcal{X}}$ admits a basis of theta functions $\{\vartheta_{p}\}_{p\in C(P)_{\mathbb{Z}}}$, indexed by the integral points of the tropicalization of $\mathcal{X}$ given by the cone $C(P)$ over $P$. It follows by Lemma 6.22 that the theta functions $\vartheta_{p}$ corresponding to the interior points of, for $p\in C(P)_{\mathbb{Z}}\setminus\partial C(P)_{\mathbb{Z}}$ define an ideal of the mirror algebra $R_{\mathcal{X}}$. We prove that this ideal defines an anticanonical divisor $\mathcal{D}_{\mathcal{X}}$ in $\mathcal{Y}_{\mathcal{X}}$. On the other hand, the intrinsic mirror family is naturally polarized by the line bundle $\mathcal{L}_{\mathcal{X}}=\mathcal{O}_{\mathcal{Y}_{\mathcal{X}}}(1)$ with a natural basis of sections $\vartheta_{p}$, for $p\in P_{\mathbb{Z}}$, and so we have natural “theta divisor” $\mathcal{C}_{\mathcal{X}}=\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\}\in|\mathcal{L}_{\mathcal{X}}|$. Hence, running the intrinsic mirror construction for every projective crepant resolution $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ of the affinization $\overline{\mathcal{X}}$ of $\mathcal{X}$, we obtain pairs $\left(\mathcal{Y}_{\mathcal{X}^{\prime}},D_{\mathcal{X}^{\prime}}+\epsilon\,\mathcal{C}_{\mathcal{X}^{\prime}}\right)$ over the affine toric varieties $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}})]$. In §6 we show that these families glue to a family over the movable secondary fan and in Theorem 6.28 we show that this family extends uniquely to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to S^{\mathrm{sec}}_{\mathcal{X}}$. Theorems 7.13 and 8.10 then prove:
@@
The intrinsic mirror family $\left(\mathcal{Y},D+\epsilon\,\mathcal{C}\right)$ to a quasi-projective open Kulikov degeneration $\mathcal{X}\to\Delta$ is a family of KSBA stable log Calabi–Yau surfaces over the toric variety $S^{\mathrm{sec}}_{\mathcal{X}}$. Moreover, when $\mathcal{X}=\mathcal{X}_{\mathscr{P}}$ is the semistable mirror of a polarized log Calabi–Yau surface $(Y,D,L)$, the general fiber of this intrinsic mirror family is deformation equivalent to $(Y,D,L)$.
To prove that the fibers $(Y_{t},D_{t}+\epsilon\,C_{t})$ of the intrinsic mirror family are slc, we first show in Theorem 7.7 that over the dense torus of the toric variety $S_{\mathcal{X}}^{\mathrm{sec}}$, the surface $Y_{t}$ is irreducible, and then it is enough to show that $C_{t}$ does not pass through the 00-strata of $D_{t}$. This follows from the description in §6.4.2 of the restriction of the theta functions to $D_{t}$. For an arbitrary fiber $(Y_{t},D_{t}+\epsilon\,C_{t})$, with irreducible components $(V_{i},D_{i}+\epsilon\,C_{i})$, one reduces using Theorem 7.12 the statement to a single pair $(V_{i},D_{i}+\epsilon\,C_{i})$ by the naturalness of the intrinsic mirror symmetry construction and the previous case.
To study the general fiber of the intrinsic mirror family to the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we show that, up to a base change of $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, a particular one-parameter sub-family of the intrinsic mirror family is a toric degeneration in the sense of Gross–Siebert [46]. This allows us to identify the general fiber topologically, and then as a log Calabi–Yau surface using a result of Friedman [34]. Then, by the universal property of the moduli space $\mathcal{M}_{(Y,D,L)}$, we obtain a map $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow\mathcal{M}_{(Y,D,L)}$. To conclude our main result, in Theorem 8.11 we show that this map is finite and surjective. The finiteness follows by proving that the intrinsic mirror family is non-constant in restriction to the one-dimensional strata of $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$, which suffices by [49, Lemma 8.7]. Finally, the surjectivity is proved using a period calculation, as in [70, 82].

### 30 Mar 2024 v2

@@ Abstract.
We prove that every irreducible component of the coarse Kollár-Shepherd-Barron and Alexeev (KSBA) moduli space of stable log Calabi–Yau surfaces is a finite quotient of a projective toric variety. This verifies a conjecture of Hacking–Keel–Yu. The proof combines tools from log smooth deformation theory, the minimal model program, punctured log Gromov–Witten theory and mirror symmetry.
@@ 1.1. Context and overview
The higher dimensional analogue of a stable curve, introduced by Kollár–Shepherd-Barron [68] and Alexeev [5], is a stable pair $(Y,B)$, consisting of a projective variety $Y$ and a $\mathbb{Q}$-divisor $B$ such that $(Y,B)$ has semi-log-canonical (slc) singularities, and satisfying the stability condition requiring $K_{Y}+B$ to be $\mathbb{Q}$-Cartier and ample – see [66] for a comprehensive treatment of the KSBA theory. The geometry of the KSBA moduli space of stable pairs has been investigated so far in a handful of cases, including toric and abelian varieties [6], K3 surfaces [7, 8, 10, 72], elliptic surfaces [20, 21], and for some surfaces of general type [9]. Most relevant to the present paper, Alexeev showed in [6] that for toric varieties the normalization of the coarse KSBA moduli space is toric, with associated fan the secondary fan defined as in [37].
More generally, for maximal log Calabi–Yau varieties, it has been conjectured by Keel [63] around 2012 that the coarse KSBA moduli space shall still admit a finite cover by a toric variety. Furthermore, Hacking–Keel–Yu conjectured that such a toric variety should be defined by a generalized version of the “secondary fan” coming from mirror symmetry [50]. We prove this conjecture in dimension two, that is, for all polarized maximal log Calabi–Yau surfaces $(Y,D,L)$ consisting of a projective surface $Y$ over $\mathbb{C}$, with a reduced, anticanonical divisor $D\subset Y$, such that $(Y,D)$ has at worst log canonical singularities, and an ample line bundle $L$ on $Y$. Furthermore, $(Y,D)$ is required to be maximal, in the sense that $D$ is non-empty and singular.
The KSBA moduli space for $(Y,D,L)$, denoted by $\mathcal{M}_{(Y,D,L)}$, is the closure in the moduli space of stable pairs of the locus of stable pairs deformation equivalent to $(Y,D+\epsilon\,C)$, where $C\in|L|$ and $10<\epsilon<\!\!<$ 1. It is a proper Deligne–Mumford stack, with a projective coarse moduli space $M_{(Y,D,L)}$. To investigate the geometry of $M_{(Y,D,L)}$ via mirror symmetry, we construct a “semistable mirror” to a maximal degeneration of $(Y,D,L)$. Such a degeneration is determined by a choice of a polyhedral decomposition $\mathscr{P}$ on a “Symington polytope” $P$ associated to $(Y,D,L)$. To obtain the semistable mirror, we first build a normal crossing surface with dual intersection complex $(P,\mathscr{P})$. The semistable mirror is then obtained as a formal smoothing of this normal crossing surface. We show that such a smoothing is an open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\to\Delta=\mathrm{Spf}\,\mathbb{C}\llbracket t\rrbracket$, given by a projective crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}_{\mathscr{P}}$ of a formal affine canonical 3-fold singularity $\overline{\mathcal{X}}_{\mathscr{P}}$. We then define the “secondary fan” $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, as a coarsening of the Mori fan for the semistable mirror to $(Y,D,L)$. Our main result, Theorem 8.11, states:
@@
The coarse KSBA moduli space $M_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface $(Y,D,L)$ admits a finite cover
$S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\longrightarrow M_{(Y,D,L)}\,,$
@@ 1.2. Motivation from mirror symmetry
Given a stable log Calabi–Yau surface $(Y,D+\epsilon\,C)$ with an ample line bundle $L$, let $U:=Y\setminus D$, and $s$ be the section of $L|_{U}$ defining $C|_{U}$. One can naturally associate to $(Y,D+\epsilon\,C)$ the non-compact Calabi–Yau 3-fold $Z$ given by the hypersurface with equation $xy=s$ in the total space of the rank two vector bundle $L|_{U}\oplus\mathcal{O}_{U}$ over $U$, where $x$ and $y$ are linear coordinates along the fibers. Said differently, $Z$ is the affine conic bundle over the non-compact Calabi–Yau surface $U$ branched over the curve $C|_{U}$. Heuristically, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ can be viewed as a compactification of a “moduli space of complex structures” on $Z$. Mirror symmetry predicts an identification of $\mathcal{M}_{(Y,D,L)}$ with the “stringy extended Kähler moduli space” of a mirror Calabi–Yau 3-fold $\mathcal{X}$ to $Z$ [77]. More precisely, under mirror symmetry different maximal degenerations of $Z$ should correspond to different mirrors related by birational transformations. In this paper, working directly with $(Y,D,L)$ instead of $Z$, we provide a precise realization of this heuristic picture, where a mathematically rigorous definition of the “stringy extended Kähler moduli space of $\mathcal{X}_{\mathscr{P}}$” is given by the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$.
@@
It is worth noting that in the situation when $Y$ is a toric variety, and $D$ is the toric boundary divisor, mirror symmetry is entirely combinatorial and it essentially reduces to a duality between fans and momentum polytopes. Let $\overline{\mathcal{X}}$ be the singular toric variety with associated fan $C(P)$, the cone over the momentum polytope $P$ of $(Y,D,L)$. Choosing a maximal regular subdivision $\mathscr{P}$ and refining $C(P)$ by adding cones over $\mathscr{P}$, one obtains a toric crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}$, the semistable mirror family to $(Y,D,L)$. By standard toric geometry, the Mori fan of this mirror family to $(Y,D,L)$ coincides with the secondary fan of $(Y,D,L)$ [25]. So, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface can be understood combinatorially, either in terms of the secondary fan, or of the Mori fan of the mirror [6, 98]. The existence of a dense torus in $\mathcal{M}_{(Y,D,L)}$ follows in this case from the existence of the basis of $H^{0}(Y,L)$ given by the monomials $z^{m}$ on $(\mathbb{C}^{\star})^{2}=Y\setminus D$ indexed by the integral points $m\in P_{\mathbb{Z}}:=P\cap\mathbb{Z}^{2}$. The key feature of this basis is that, for a curve $C$ defined by an equation $\sum_{m\in P_{\mathbb{Z}}}a_{m}z^{m}=0$ with $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$, the pair $(Y,D+\epsilon\,C)$ is KSBA-stable. Indeed, since for any vertex $m^{\prime}$ of $P$, having $a_{m^{\prime}}\in\mathbb{C}^{\star}$ ensures that $C$ does not intersect the singular point of $D$ corresponding to $m^{\prime}$, it follows that $(Y,D+\epsilon\,C)$ has only slc singularities if $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$ [6].
@@ 1.3. Outline of the paper
In §2, we investigate maximal degenerations of polarized log Calabi–Yau surfaces combinatorially, using real affine geometry. To do so, we consider a “nice” toric model of a polarized log Calabi–Yau $(Y,D,L)$, as in previous work of Engel–Friedman [31], which allows us to systematically describe an associated “Symington polytope” $P$ to $(Y,D,L)$ given by an integral affine manifold with singularities [91]. We define the notion of a good polyhedral decomposition on $P$, and we prove the existence of such decompositions in Theorem 2.14. A good polyhedral decomposition $\mathscr{P}$ on $P$ defines in particular a maximal degeneration of $(Y,D,L)$, whose central fiber has “intersection complex” $(P,\mathscr{P})$, as explained in §2.4.
To construct a “mirror” to such a maximal degeneration of $(Y,D,L)$, we first construct a normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ with “dual intersection complex” $(P,\mathscr{P})$ in §4.2. The irreducible components of $\mathcal{X}_{\mathscr{P},0}$ are obtained by $\mathbb{Q}$-Gorenstein smoothings of toric surfaces with quotient Wahl singularities, using Lemmas 4.1 and 4.2 which show that there are no local to global obstructions for such smoothings. Finally, we appropriately glue these irreducible components to obtain the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$, which has features similar to the central fiber of a Type III Kulikov degeneration in the context of K3 surfaces [35]. A key difference in our context is that the dual intersection complex of $\mathcal{X}_{\mathscr{P},0}$ is a disk, rather than a sphere. So, in §3 we introduce the general notion of an open Kulikov surface, as a normal crossing surface $\mathcal{X}_{0}$ with dual intersection complex a disk, and satisfying some additional properties – see Definition 3.7. After investigating properties of such surfaces, as well as of their affinizations, we show in §4 that for an appropriate gluing the surface $\mathcal{X}_{\mathscr{P},0}$ associated is a $d$-semistable open Kulikov surface. Our key results in Theorems 3.26 and 3.29 prove the following result for a general open Kulikov surface $\mathcal{X}_{0}$:
@@
We refer to such a smoothing $\mathcal{X}\to\Delta$ as an open Kulikov degeneration. By the Minimal Model Program, adapted to the formal set up [75], we conclude in §3.6 that such a Kulikov degeneration is a relative Mori dream space. Hence, we obtain a complete toric fan, the “MoriFan”, in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})\otimes\mathbb{R}$. Then, following [50], we define the “secondary fan” $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$ as a “coarsening” of this Mori fan, by identifying neighboring cones connected by “M1 flops” – for details see §6.4, and denote by $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ the corresponding toric variety.
For a polarized log Calabi–Yau surface $(Y,D,L)$ with a choice of $(P,\mathscr{P})$, we refer to the open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ obtained by smoothing $\mathcal{X}_{\mathscr{P},0}$ as the “semistable mirror” of $(Y,D,L)$. We then investigate the “double mirror”, obtained by applying a version of the intrinsic mirror construction of Gross–Siebert [48] to the semistable mirror $\mathcal{X}_{\mathscr{P}}\to\Delta$. More generally, we construct an intrinsic mirror for any quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$. One subtlety in our setup is that $\mathcal{X}\rightarrow\Delta$ is only quasi-projective, while the intrinsic mirror construction of [48] applies to projective degenerations. Therefore, we first prove in Theorem 5.2 that the moduli spaces of punctured log Gromov–Witten invariants of an open Kulikov surface $\mathcal{X}_{0}$, defining the structure constants in the intrinsic mirror construction, are proper. The proof crucially uses the fact that we have a projective crepant resolution $\mathcal{X}\to\overline{\mathcal{X}}$, where $\overline{\mathcal{X}}$ is affine. Hence, analogously as in [48], we define a commutative and associative graded mirror algebra $\mathcal{R}_{\mathcal{X}}$ – see Theorem 5.5. Thus, taking the Proj of this algebra, we obtain an intrinsic mirror family to $\mathcal{X}\to\Delta$ a priori defined over the formal scheme $\mathrm{Spf}\,\mathbb{C}\llbracket NE(\mathcal{X}/\overline{\mathcal{X}})\rrbracket$, as in [48], where $NE(\mathcal{X}/\overline{\mathcal{X}})$ stands for the monoid of integral points in the cone of relative effective curve classes of $\mathcal{X}\rightarrow\overline{\mathcal{X}}$.
@@
By construction [48], the intrinsic mirror algebra $\mathcal{R}_{\mathcal{X}}$ admits a basis of theta functions $\{\vartheta_{p}\}_{p\in C(P)_{\mathbb{Z}}}$, indexed by the integral points of the tropicalization of $\mathcal{X}$ given by the cone $C(P)$ over $P$. It follows by Lemma 6.22 that the theta functions $\vartheta_{p}$ corresponding to the interior integral points of $C(P)$ define an ideal of the mirror algebra $R_{\mathcal{X}}$. We prove that this ideal defines an anticanonical divisor $\mathcal{D}_{\mathcal{X}}$ in $\mathcal{Y}_{\mathcal{X}}$. On the other hand, the intrinsic mirror family is naturally polarized by the line bundle $\mathcal{L}_{\mathcal{X}}=\mathcal{O}_{\mathcal{Y}_{\mathcal{X}}}(1)$ with a natural basis of sections $\vartheta_{p}$, for $p\in P_{\mathbb{Z}}$, and so we have a natural “theta divisor” $\mathcal{C}_{\mathcal{X}}:=\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\}\in|\mathcal{L}_{\mathcal{X}}|$. Hence, running the intrinsic mirror construction for every projective crepant resolution $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ of the affinization $\overline{\mathcal{X}}$ of $\mathcal{X}$, we obtain pairs $\left(\mathcal{Y}_{\mathcal{X}^{\prime}},D_{\mathcal{X}^{\prime}}+\epsilon\,\mathcal{C}_{\mathcal{X}^{\prime}}\right)$ over the affine toric varieties $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}})]$. In §6 we show that these families glue to a family over the movable secondary fan and in Theorem 6.28 we show that this family extends uniquely to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to\mathcal{S}^{sec}_{\mathcal{X}}$, where $\mathcal{S}^{sec}_{\mathcal{X}}$ is a toric Deligne-Mumford stack defined by a stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ with underlying fan the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$. Theorems 7.13 and 8.10 then prove:
@@
The intrinsic mirror family $\left(\mathcal{Y},D+\epsilon\,\mathcal{C}\right)$ to a quasi-projective open Kulikov degeneration $\mathcal{X}\to\Delta$ is a family of KSBA stable log Calabi–Yau surfaces over the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. Moreover, when $\mathcal{X}=\mathcal{X}_{\mathscr{P}}$ is the semistable mirror of a polarized log Calabi–Yau surface $(Y,D,L)$, the general fiber of this intrinsic mirror family is deformation equivalent to $(Y,D,L)$.
To prove that the fibers $(Y_{t},D_{t}+\epsilon\,C_{t})$ of the intrinsic mirror family are slc, we first show in Theorem 7.7 that over the dense torus of the toric stack $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$, the surface $Y_{t}$ is irreducible, and then it is enough to show that $C_{t}$ does not pass through the 00-strata of $D_{t}$. This follows from the description in §6.4.2 of the restriction of the theta functions to $D_{t}$. For an arbitrary fiber $(Y_{t},D_{t}+\epsilon\,C_{t})$, with irreducible components $(V_{i},D_{i}+\epsilon\,C_{i})$, we show using Theorem 7.12 that it suffices to prove that the pairs $(V_{i},D_{i}+\epsilon\,C_{i})$ are slc. By the naturalness of the intrinsic mirror symmetry, this follows from Theorem 7.7.
To study the general fiber of the intrinsic mirror family to the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we show that, up to a base change of $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, a particular one-parameter sub-family of the intrinsic mirror family is a toric degeneration in the sense of Gross–Siebert [47]. This allows us to identify the general fiber topologically, and then as a log Calabi–Yau surface using a result of Friedman [34]. Then, by the universal property of the moduli space $\mathcal{M}_{(Y,D,L)}$, we obtain a map of stacks $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow\mathcal{M}_{(Y,D,L)}$, inducing a map $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow M_{(Y,D,L)}$ at the level of coarse moduli spaces. To conclude our main result, in Theorem 8.11 we show that this map is finite and surjective. The finiteness follows by proving that the intrinsic mirror family is non-constant in restriction to the one-dimensional strata of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$, which suffices by [50, Lemma 8.7]. Finally, the surjectivity is proved using a period calculation, as in [71, 83].